# Assessing the Macro and Micro Effects of Random Seeds on Fine-Tuning Large Language Models

### v1

@@ Abstract
The impact of random seeds in fine-tuning large language models (LLMs) has been largely overlooked despite its potential influence on model performance. In this study, we systematically evaluate the effects of random seeds on LLMs using the GLUE and SuperGLUE benchmarks. We analyze the macro-level impact through traditional metrics like accuracy and F1, calculating their mean and variance to quantify performance fluctuations. To capture the micro-level effects, we introduce a novel metric, consistency, measuring the stability of individual predictions across runs. Our experiments reveal significant variance at both macro and micro levels, underscoring the need for careful consideration of random seeds in fine-tuning and evaluation.
@@
Hao Zhou1, Guergana Savova2, Lijing Wang1 1New Jersey Institute of Technology 2Boston Children’s Hospital and Harvard Medical School
@@ 1 Introduction
The impact of random seeds in neural network training has long been recognized across various domains, such as general machine learning classification and regression tasks Ganesh et al. (2023); Madhyastha and Jain (2019), computer vision Picard (2021); Åkesson et al. (2024), natural language processing (NLP)Bethard (2022); Lucic et al. (2022). Random seeds influence initialization and training dynamics, introducing variability in model outcomes that can lead to significant fluctuations in performance Bengio (2012).
In the field of NLP, large language models (LLMs) have achieved remarkable success across a wide range of NLP tasks, setting new state-of-the-art results on benchmarks like GLUE and SuperGLUE. These benchmarks have became the de facto standard for evaluating the capabilities of LLMs in understanding, reasoning, and generating natural language. Despite the success of LLMs, the pretrained transformer architectures, such as BERT Devlin et al. (2019) and RoBERTa Liu (2019), have been found to be particularly sensitive to random seeds Risch and Krestel (2020); Dodge et al. (2020); Mosbach et al. (2021). This sensitivity can result in substantial performance variations, complicating the interpretation of experimental results and undermining confidence in benchmarking or state-of-the-art outcomes. A recent analysis of 85 papers from the ACL Anthology Bethard (2022) revealed risky practices in the use of random seeds: over 50% of the papers exhibited potential misuse, with 24 using a single fixed random seed. This highlights that the influence of random seeds on LLM performance is still an underexplored area.
@@
Existing studies examining the impact of random seeds Ganesh et al. (2023); Madhyastha and Jain (2019); Picard (2021) typically evaluate performance variations by measuring the variance of standard metrics, such as accuracy and F1 score for classification tasks, or Pearson correlation for regression tasks, across multiple seeds. These evaluations focus on the macro-level agreement of model performance across the entire test set, offering insights into overall variability. However, they overlook the micro-level impact of how individual test points are influenced by random seed variations. As shown in Figure 1, model performance is robust to random seeds 42 and 52 at the macro level (both achieve 60% accuracy) but lacks consistency at the micro level (only 20% overlapping predictions). This micro-level inconsistency can have severe consequences in real-world applications, especially in fields where model predictions are highly sensitive to individual test points, such as medical diagnosis and autonomous driving. Understanding this micro-level effect is crucial for assessing model robustness at the level of individual predictions, ensuring that specific test samples are not inconsistently misclassified or predicted due to seed-induced variations. Additionally, it helps pinpoint specific areas where models may exhibit significant instability, such as consistently misclassifying certain types of data points or showing highly variable predictions for similar inputs. Recognizing these areas of instability can guide targeted improvements in both model design and evaluation practices, ensuring that assessments account for seed-induced variations in performance.
Major contributions: To address these gaps, in this work, (1) we analyze the impact of random seeds on pretrained LLMs using the GLUE and SuperGLUE benchmarks, covering both macro and micro-level variability; (2) We introduce a novel consistency metric to assess prediction stability on individual test points, capturing the micro-level effects of random seeds; (3) Our extensive experiments reveal significant variability in both standard and consistency metrics, underscoring the need to consider seed-induced variations in fine-tuning and evaluation, and incorporate random seed sensitivity into benchmarking and reporting for more reliable and reproducible results.
@@ 2 Macro Metric: Variance
To measure the macro-level impact of random seeds on LLM performance, we calculate the variance of a standard metric across multiple seeds. Let $[\zeta_{1},\cdots,\zeta_{S}]$ represent the values of a model performance metric for LLMs fine-tuned with $S$ random seeds, the variance is calculated by:
@@
where $\bar{\zeta}=\frac{1}{S}\sum_{i=1}^{S}\zeta_{i}$. $\zeta$ can be any standard metrics, such as F1 score for classification tasks or Pearson correlation for regression tasks. A smaller VAR indicates less variation in macro-level performance.
@@ 3 Micro Metric: Consistency
’Consistency’ can have varying definitions across domains. Building on prior work, Wang et al. (2020) formally defined the consistency of a deep learning model as its ability to produce consistent predictions for the same input when periodically retrained with streaming data in deployment settings. Extending this idea, we define the consistency of an LLM as its ability to generate consistent predictions for the same input across models fine-tuned with different hyperparameter settings, with correct-consistency further specifying its ability to make consistent correct predictions in this context.
More specifically, consider two LLMs $A$ and $B$, given a dataset $\mathcal{D}={d_{1},\cdots,d_{N}}$ of $N$ data points, $y_{i}^{A}$ and $y_{i}^{B}$ are the prediction of $A$,$B$ for a data point $d_{i}$ with ground truth $r_{i}$. For classification tasks, we calculate the consistency as follows:
[…]
where $1_{A,B}(\cdot)$ is the indicator function that equals 1 if $y_{t}^{A}=y_{t}^{B}$, otherwise 0. And the correct-consistency is calculated by:
[…]
where $1_{A,B,r}(\cdot)$ is the indicator function that equals 1 if $y_{t}^{A}=y_{t}^{B}=r_{t}$, otherwise 0.
For regression tasks, we calculate the consistency as the Pearson correlation between two sets of predictions $[y_{1}^{A},...,y_{N}^{A}]$ and $[y_{1}^{B},...,y_{N}^{B}]$:
[…]
CON and ACC-CON range from [0, 1], while CON-PEAR ranges from [-1, 1], where higher values indicate smaller variations in micro-level predictions. While consistency metrics can generally be used for quantifying the agreement of individual predictions from any two LLMs with different architectures, hyperparameters, or training settings, in our study, they are specifically used to serve as metrics to evaluate the micro-level impact of random seeds on the same pretrained LLM.
@@ 4 Experimental Setup
In this section, we will show the details of our experiment settings. More setting details can be found in Appendix.
4.1 Benchmarks
In this study, we conduct experiments on a range of NLP tasks including COLA (Matthews Correlation Coefficient), SST2 (Accuracy), MRPC (Accuracy), STSB (Pearson correlations), QQP (Accuracy), QNLI (Accuracy), and RTE (Accuracy) from GLUE Wang et al. (2018) benchmark; RTE (Accuracy), CB (Accuracy), WiC (Accuracy), BoolQ (Accuracy), MultiRC (Accuracy), and COPA (Accuracy) from SuperGLUE Wang et al. (2019) benchmark. In our paper, we use RTEG to denote RTE task from GLUE and RTES for SuperGLUE.
To ensure our experiments are conducted with proper settings and are reproducible, we followed the configurations and replicated the state-of-the-art (SOTA) scores reported in Liu (2019). We choose RoBERTa-large as the pretrained LLM. We were unable to replicate the representation (special token extractions) and model settings (unpublished pretrained model) for the WSC and MNLI tasks, so they are omitted from the experiment.
@@ 4.2 Settings
Our experiments were implemented using Hugging Face Transformers (v4.30.0) and PyTorch (v2.0), conducted on NVIDIA A100 GPUs with 40GB of memory across two GPUs. With limited computational resources, we perform fully fine-tuning for each task with five random seeds: 42, 52, 62, 72, 82 (i.e., $S=5$), which are randomly chosen except 42 (see Section 5.1). To calculate CON, ACC-CON, and CON-PEAR, we compute the metric for each of the 10 random seed pairs (e.g., 42 and 52) and then take the average of these 10 values as the final consistency score. We used the run_glue.py PyTorch script for fine-tuning, and default settings were applied unless otherwise specified. Although differences in the fine-tuning script and missing settings from the original authors prevented us from reproducing the exact SOTA scores, our results are close to the reported SOTA. A comparison of our implementation with the reference SOTA scores and detailed data and learning settings are provided in Appendix Table 2, Table 3, and Table 4.
SuperGLUE BoolQ CB RTES MultiRC WiC COPA
ACC 85.05 98.8 69.6 79.01 68.4 73.2
VAR 0.24 1.1 18.22 12.21 2.83 12.83
CON 94.78 91.61 71.84 76.16 79.09 67.6
ACC-CON 82.44 90.18 56.61 67.10 57.95 57
GLUE MRPC QNLI QQP SST2 RTEG COLA STSB
ACC 90.34 94.53 92.02 95.73 84.04 64.51 92.19
VAR 0.93 0.16 0.07 0.71 0.47 0.64 0.30
CON 92.21 96.84 96.00 98.10 92.78 94.23 0.9853
ACC-CON 52.33 92.95 90.03 94.58 80.43 82.52 -
Table 1: Macro- and micro-impact of five random seeds. ACC is the average of five accuracies (SOTA metrics). VAR is the variance of Accuracy calculated using Equation 1. CON and ACC-CON are the average of 10 values, each derived from Equation 2, 3, or 4. For STSB, we put CON-PEAR value in CON row for concise format. ACC, CON, and ACC-CON are expressed as percentages.
@@ 5 Results and Discussion
In this section, we will show experimental results at both macro- and micro-level and discuss key findings.
@@ 5.1 Macro impact
Table 1 presents the averaged accuracy (ACC) and variance (VAR) for GLUE and SuperGLUE tasks across five random seeds. Significant variance in macro-level performance is observed in many tasks, reflecting sensitivity to random seed selection. Tasks like RTES (VAR = 18.22), COPA (VAR = 12.83), and MultiRC (VAR = 12.21) exhibit high variability in ACC, highlighting the need for robust evaluation methods and stability-enhancing techniques, such as more robust optimization methods, better hyperparameter tuning, or ensembling across multiple seeds. In contrast, tasks like QQP (VAR = 0.07) and QNLI (VAR = 0.16) show much greater stability, likely due to their inherent properties such as larger datasets or simpler decision boundaries. High variability undermines the reliability of single-seed evaluations, emphasizing the importance of averaging results and addressing task-specific challenges to improve model robustness.
There is a "common belief" in the machine learning community that random seed 42 may outperform others. To investigate whether a specific random seed consistently leads to better results across different models or tasks, in Figure 2 we present a heatmap of normalized ACC for each task across five random seeds. There is no significant difference in color distribution between each row, indicating that no discernible pattern or evidence supporting the existence of a universally superior random seed.
Figure 2: A heatmap of normalized ACC across tasks and five random seeds, with a darker color representing a better accuracy.
@@ 5.2 Micro impact
Table 1 reports consistency (CON) and correct-consistency (ACC-CON) for GLUE and SuperGLUE tasks across five random seeds. High CON values in tasks like SST2 (98.1%), QNLI (96.84%), and QQP (96%) indicate stable predictions, while lower values for RTES (71.84%) and COPA (67.7%) highlight their sensitivity to random seeds, potentially due to smaller training sizes or task complexity. High ACC-CON in SST2 (94.58%) and QNLI (92.95%) suggest stable correct predictions, whereas low ACC-CON in RTES (56.6%) and MRPC (52.33%) reveal that consistent predictions are not always accurate, emphasizing the need to evaluate both stability and correctness. Additionally, MRPC’s low VAR (0.93) value demonstrates that similar macro-level accuracy does not necessarily imply true reproducibility, underscoring the importance of micro-level analysis beyond macro-level metrics.
Identifying robust data points—those consistently predicted correctly through micro-level analysis—and leveraging them to enhance data collection, preprocessing, prompt engineering, or synthetic data generation offer a potential solution for mitigating seed-induced variability and improving LLM robustness.
5.3 Training size impact
@@
Training size significantly influences a model’s predictive performance, with larger datasets generally improving accuracy, though this is not guaranteed due to factors like task complexity and label noise Shahinfar et al. (2020); Althnian et al. (2021); Bailly et al. (2022). Will increasing training data size improve variance and consistency in general? To answer the question, we show Pearson correlation analysis between training size, variance, and consistency in Figure 3. It reveals a weak negative correlation (-0.25) between training size and VAR, indicating that smaller datasets tends to increase performance variance, as seen in RTES (highest VAR of 18.22 with relatively small training size). However, the effect is not pronounced or consistent across all tasks, as MultiRC and WiC exhibit high VAR despite a relatively large dataset. A weak or moderate positive correlation is observed between training size and both CON (0.41) and ACC-CON (0.43), suggesting larger datasets generally improve consistency and prediction stability across random seeds, but with no guarantee.
Increasing training size can reduce both macro and micro variability to random seeds, but its effectiveness depends on factors like data quality, task complexity, and label noise. Alternatively, as discussed in Section 5.2, identifying robust data points and augmenting the training data with data points having similar robust patterns (either real data or generated synthetic data) provide a more targeted strategy to mitigate seed-induced variability and improve LLM robustness.
@@
Figure 3: Correlation between training size (log scale), VAR, CON, and ACC-CON. Tasks are arranged in ascending order of training size, with exact sizes detailed in Appendix 3.

### v2

@@ Abstract
The impact of random seeds in fine-tuning large language models (LLMs) has been largely overlooked despite its potential influence on model performance. In this study, we systematically evaluate the effects of random seeds on LLMs using the GLUE and SuperGLUE benchmarks. We analyze the macro impact through traditional metrics like accuracy and F1, calculating their mean and variance to quantify performance fluctuations. To capture the micro effects, we introduce a novel metric, consistency, measuring the stability of individual predictions across runs. Our experiments reveal significant variance at both macro and micro levels, underscoring the need for careful consideration of random seeds in fine-tuning and evaluation.
@@
Nghia T. Bui1, Guergana Savova2, Lijing Wang1 1New Jersey Institute of Technology {ntb23, lijing.wang}@njit.edu 2Boston Children’s Hospital and Harvard Medical School guergana.savova@childrens.harvard.edu
@@ 1 Introduction
The impact of random seeds in neural network training has long been recognized across various domains, such as general machine learning classification and regression tasks ganesh2023impact madhyastha2019model, computer vision picard2021torch aakesson2024random, natural language processing (NLP)bethard2022we,lucic-etal-2022-towards.
In the field of NLP, large language models (LLMs) have achieved state-of-the-art results on benchmarks like GLUE and SuperGLUE, which are now standard for evaluating language understanding and reasoning. However, pretrained transformers such as BERT devlin-etal-2019-bert and RoBERTa liu2019roberta are highly sensitive to random seeds risch-krestel-2020-bagging; dodge2020fine; mosbach2020stability, often leading to significant performance variation that complicates experimental interpretation and benchmarking. While other sources of randomness, such as prompt formatting he2024does, in-context example selection gupta2023coverage, and how learnable weights are initialized hayou2024impact, have also been explored, seed variation remains a fundamental and underaddressed issue. A recent analysis of 85 papers from the ACL Anthology bethard2022we revealed risky practices in the use of random seeds: over 50% of the papers exhibited potential misuse, with 24 using a single fixed random seed. This highlights that random seeds sensitivity in LLM fine-tuning remains insufficiently understood, motivating the need for more systematic investigation.
@@
Existing studies examining the impact of random seeds ganesh2023impact; madhyastha2019model; picard2021torch typically evaluate performance variations by measuring the variance of standard metrics, such as accuracy score for classification tasks, or MAE (mean absolute error) for regression tasks, across multiple seeds. These evaluations focus on the macro agreement of model performance across the entire test set, offering insights into overall variability. However, they overlook the micro impact of how individual test points are influenced by seed-induced variability. As shown in Figure 1, model performance is robust to random seed 42 and 52 at the macro level (both achieve 60% accuracy) but lacks consistency at the micro level (only 20% overlapping predictions). This micro inconsistency can have severe consequences in real-world applications, especially in fields where model predictions are highly sensitive to individual test points, such as medical diagnosis and autonomous driving. Understanding this micro effect is crucial for assessing model robustness at the level of individual predictions, ensuring that specific test samples are not inconsistently predicted due to seed-induced variability. Additionally, it helps pinpoint specific areas where models may exhibit significant instability, such as consistently misclassifying certain types of data points or showing highly variable predictions for similar inputs. Recognizing these areas of instability can guide targeted improvements in both model design and evaluation practices, ensuring that assessments account for seed-induced variability in performance.
Major contributions: To address these gaps, in this work, (1) we analyze the impact of random seeds on pretrained LLMs using the GLUE and SuperGLUE benchmarks, covering both macro and micro variability; (2) We introduce a novel consistency metric to assess prediction stability on individual test points, capturing the micro effects of random seeds; (3) Our extensive experiments reveal significant variability in both standard and consistency metrics, underscoring the need to consider seed-induced variability in fine-tuning and evaluation, and incorporate random seeds sensitivity into benchmarking and reporting for more reliable and reproducible results.
@@ 2 Macro Metric: Variance
To measure the macro impact of random seeds on LLM performance, we calculate the variance of a standard evaluation metric across multiple seeds. Let $\zeta_{1},\cdots,\zeta_{S}]$ represent the values of model performance for an LLM fine-tuned with $S$ random seeds, the variance is calculated by:
@@
where $\bar{\zeta}=\frac{1}{S}\sum_{i=1}^{S}\zeta_{i}$. $\zeta$ can be any standard metrics, such as accuracy score for classification tasks or MAE for regression tasks. A smaller VAR indicates less variation in macro performance.
@@ 3 Micro Metric: Consistency
’Consistency’ can have varying definitions across domains. Building on prior work, wang2020wisdom formally defined the consistency of a deep learning model as its ability to produce consistent predictions for the same input when periodically retrained with streaming data in deployment settings. Extending this idea, we define the consistency of an LLM as its ability to generate consistent predictions for the same input across models fine-tuned with different hyperparameter settings, with correct-consistency further specifying its ability to make consistent correct predictions in this context.
More specifically, consider two LLMs $A$ and $B$, given a dataset $\mathcal{D}={d_{1},\cdots,d_{N}}$ of $N$ data points, $y_{i}^{A}$ and $y_{i}^{B}$ are the prediction of AA,$B$ for a data point $d_{i}$ with ground truth $r_{i}$. The consistency of $A$ and $B$ can be calculated as follows:
$\text{CON:}\frac{1}{N}\sum_{t=1}^{N}{\pi_{A,B}(t)}$ (2)
where $\pi_{A,B}(\cdot)$ is the scoring function that quantifies the alignment between predictions $y_{t}^{A}$ and $y_{t}^{B}$, with higher values indicating smaller variations in micro-level predictions; it can be either binary (e.g., 1 for match, 0 for mismatch) or probabilistic based on different NLP tasks. And the correct-consistency is calculated by:
$\text{CCON:}\frac{1}{N}\sum_{t=1}^{N}{\pi_{A,B,r}(t)}$ (3)
where $\pi_{A,B,r}(\cdot)$ is the scoring function that quantifies the alignment between predictions $y_{t}^{A}$, $y_{t}^{B}$, and ground truth $r_{t}$.
In this paper, we focus solely on classification tasks because the GLUE and SuperGLUE benchmarks primarily consist of classification problems. The scoring function $\pi(\cdot)$ is defined as an indicator function that equals 1 if $y_{t}^{A}=y_{t}^{B}(=r_{t})$, otherwise 0. We summarize the standard metric $\zeta$ and the possible corresponding scoring functions $\pi$ for various NLP tasks in Table 2 in the Appendix A.1. We hope this summary offers useful context for interpreting our results and supports future extensions of our evaluation to a broader range of NLP tasks.
While consistency metrics can generally be used for quantifying the agreement of individual predictions from any two LLMs with different architectures, hyperparameters, or training settings, in our study, they are specifically used to serve as metrics to evaluate the micro impact of random seeds on the same pretrained LLM.
RoBERTa-large Llama3.2-3B
GLUE MRPC QNLI QQP SST2 RTEG CoLA GLUE MRPC QNLI QQP SST2 RTEG CoLA
$\zeta$ 90.34 94.00 92.00 95.59 85.02 65.61 $\zeta$ 84.02 94.20 89.37 96.78 85.92 61.33
VAR 0.89 0.38 0.06 0.55 1.48 1.32 VAR 0.56 0.14 0.09 0.27 3.34 0.88
CON 92.89 95.64 95.57 96.83 91.09 93.95 CON 90.34 96.85 96.63 98.32 88.80 94.86
CCON 86.79 91.80 89.79 94.01 80.55 82.76 CCON 79.19 92.63 87.69 95.95 80.32 81.36
SuperGLUE BoolQ CB RTES MultiRC WiC COPA SuperGLUE BoolQ CB RTES MultiRC WiC COPA
$\zeta$ 83.05 94.64 72.89 76.46 68.57 73.00 $\zeta$ 72.49 73.92 68.66 80.14 68.71 84.20
VAR 7.35 3.95 16.67 13.30 2.77 16.83 VAR 1.62 2.03 2.77 0.57 1.58 5.17
CON 86.79 91.50 70.05 74.63 80.36 64.88 CON 81.10 89.82 71.69 86.73 85.86 86.8
CCON 76.37 89.84 64.58 63.77 58.75 55.44 CCON 63.09 70.35 54.51 73.50 61.64 77.6
Table 1: Macro and micro impact of ten random seeds. $\zeta$ is the average of 10 values, which is MCC for CoLA and accuracy for the other tasks. VAR is the variance of $\zeta$ calculated using Equation 1. CON and CCON are the average of 45 consistency values calculated using Equation 2 and 3. $\zeta$, CON, and CCON are expressed as percentages.
@@ 4 Experimental Setup
4.1 Benchmarks and pretrained models
In this study, we conduct experiments on a range of NLP tasks including CoLA, SST2, MRPC, QQP, QNLI, and RTE from GLUE wang2018glue benchmark; RTE, CB, WiC, BoolQ, MultiRC, and COPA from SuperGLUE wang2019superglue benchmark. STSB, WSC, and MNLI tasks are omitted from our experiment. We specify the reason in Section A.4. All tasks use accuracy as the standard evaluation metric $\zeta$, except for CoLA, which uses Matthews Correlation Coefficient (MCC). In our paper, we use RTEG to denote RTE task from GLUE and RTES for SuperGLUE.
To examine the effects on various scales of LLMs, we experiment with RoBERTa-large (~350M trainable parameters), as well as a larger LLM Llama3.2-3B (~3.21B trainable parameters) using LoRA hu2022lora fine-tuning, enabling us to assess whether our findings generalize across model scales.
@@ 4.2 Settings
Our experiment is implemented using Hugging Face Transformers (v4.30.0) and PyTorch (v2.0), conducted on NVIDIA two A100 GPUs with 80GB of memory each. Based on the empirical findings in  wang-etal-2023-two; dodge2020fine; mosbach2020stability, 5–10 seeds are sufficient to estimate variance of LLMs in NLP tasks. We perform full fine-tuning for each task with ten randomly chosen seeds: 42, 52, 62, 72, 82, 92, 102, 112, 122, 132 (i.e., $S=10$ ). We calculate CON and CCON on each unique pair of seeds and report the average of 45 values as the final consistency score. Our fine-tuning process is based on the PyTorch script run_glue.py, and the best previously reported settings were applied unless otherwise specified.
To ensure proper experimental setup and reproducibility, we refer the configurations in liu2019roberta and replicate state-of-the-art (SOTA) scores reported using RoBERTa-large with full fine-tuning. A comparison of our implementation with the reference SOTA scores and detailed data and learning settings are provided in Appendix Table 3, Table 4, and Table 5. Since we could not find reference configurations or performance reports for Llama3.2-3B on the two benchmarks, we use the SOTA scores (Table 3) as a reference and conduct experiments using our own settings, detailed in Table 6.
@@ 5.1 Macro impact
Table 1 presents the averaged accuracy ($\zeta$) and variance (VAR) for all tasks across ten random seeds using two LLMs. Significant VAR is observed in many tasks using RoBERTa-large, such as RTES (16.67), COPA (16.83), and MultiRC (13.30), reflecting sensitivity to random seed selection. High variability at the macro level undermines the reliability of single-seed evaluations, emphasizing the need for robust evaluation methods and stability-enhancing techniques. In contrast, tasks like QQP (0.06), QNLI (0.38), SST2 (0.55), and MRPC (0.89) show much greater stability, likely due to their inherent properties such as larger datasets or simpler decision boundaries. This also helps explain why SuperGLUE tasks generally show higher VAR than GLUE tasks.
Compared to RoBERTa-large, Llama3.2-3B with LoRA fine-tuning exhibits significantly lower VAR across most tasks. This is likely because only a small subset of parameters (~2.3 million) is updated during LoRA fine-tuning, which constrains the variance introduced by random seeds and results in greater stability and robustness to seed-induced fluctuations. Furthermore, the decoder-only, autoregressive architecture of Llama3.2-3B may be inherently less sensitive to minor parameter perturbations during task adaptation compared to the encoder-only structure of RoBERTa, contributing to its robust performance. A deeper analysis of how these factors govern performance variance is a promising avenue for future work.
@@ 5.2 Micro impact
Table 1 reports consistency (CON) and correct-consistency (CCON) for all tasks over ten random seeds. For RoBERTa-large, high CON values in tasks like SST2 (96.83), QNLI (95.64), and QQP (95.57) indicate stable predictions, while lower values in tasks like COPA (64.88) and RTES (70.05) highlight their sensitivity to random seeds, potentially due to smaller training sizes or task complexity. Tasks with large CON–CCON gaps (e.g., WiC with a 21.61 difference and CoLA with 11.19) suggest that consistent predictions are not always accurate, emphasizing the need to evaluate both stability and correctness. Furthermore, tasks like WiC, which show low VAR alongside low CON and CCON, demonstrate that similar macro accuracy can mask underlying instability, reinforcing the importance of micro-level evaluation beyond traditional metrics. Results from Llama3.2-3B show consistent trends, albeit with task-specific variations.
Unlike the macro impact where Llama3.2-3B with LoRA fine-tuning shows significantly lower VAR than RoBERTa-large, no such trend is observed for CON and CCON. This implies that using a parameter-efficient method like LoRA with a modern LLM helps mitigate macro variance but has limited effect on micro consistency.
5.3 Discussion
Will increasing training data size improve variance and consistency in general?
@@ 5.3 Discussion
To answer the question, we present Pearson correlation analysis in Figure 2, showing the relationship between training size, variance, and consistency, with tasks sorted by increasing dataset size. It reveals a weak negative correlation (-0.3918) between training size and VAR, indicating that smaller datasets tend to increase macro variance. However, the effect is not pronounced or consistent across all tasks, as MultiRC exhibits high VAR despite a relatively large dataset. A weak or moderate positive correlation is observed between training size and both CON (0.4257) and CCON (0.4259), suggesting that larger datasets generally improve consistency and prediction stability across random seeds, but with no guarantee. Increasing training size can reduce macro and micro variability in random seeds, but its effectiveness depends on factors like data quality, task complexity, and label noise shahinfar2020many; althnian2021impact; bailly2022effects.
@@
Figure 2: Correlation between training size (log scale), VAR, CON, and CCON while using RoBERTa-large. Tasks are arranged in ascending order of training size, with exact sizes detailed in Appendix 4.
Outlook: mitigating seed-induced variability.
Our analysis reveals substantial seed-induced variability in model performance at both the macro and micro levels. The findings indicate that increasing the training size can reduce both macro and micro variability in random seeds, but it’s not a guaranteed solution, especially for complex tasks.
Prior studies have proposed several strategies to mitigate macro variability, including model ensembling risch-krestel-2020-bagging; wang-etal-2023-two; wang2020wisdom, stability-aware training dodge2020fine, and more robust evaluation protocols mosbach2020stability. Among these methods, only wang2020wisdom explicitly addresses consistency through snapshot ensembles. While this method guarantees improvement, it is computationally prohibitive for LLMs, as it requires training numerous models to form a sufficient ensemble.
Inspired by our observation that individual predictions are highly sensitive to randomness, a promising future direction is the development of novel optimization algorithms. Specifically, one could dynamically weight the loss of individual examples during gradient propagation to stabilize training and reduce both macro and micro variance, without incurring significant computational overhead. We leave the exploration of such methods for future work.